OpenStudy (baby456):

Please help for medal! Determine the slop of the line through the points (-1/2 ,-2) and (3/2 , 1)

OpenStudy (baby456):

@mathmale

OpenStudy (baby456):

@zzr0ck3r

OpenStudy (baby456):

@jdoe0001

OpenStudy (zzr0ck3r):

for points \((a,b)\) and \((c,d)\) the slope of the line going through them is given by \[m=\frac{a-c}{b-d}\]

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -\frac{1}{2}}}\quad ,&{\color{blue}{ -2}})\quad &({\color{red}{ \frac{3}{2}}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)